Circular No. 8730

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/1999 X1 = 2006 O1 (HUG-BELL) D. Tibbets and G. Hug, Eskridge, KS, report the CCD recovery of

comet P/1999 X1 at mag 18.6–18.7. Astrometry, new orbital elements (indicated correction of ΔT = −0.12 day to the prediction; cf. MPC 48383, 2006 Comet Handbook), and an ephemeris appear on MPEC 2006-O03.

2006 UT  α_{2000}  δ_{2000}
July 16.40474  4^{h}28^{m}21.^{s}65  +14^{o}24′28′′.0

SATURN XXXV (DAPHNIS) = S/2005 S 1 Further to IAUC 8471, the IAU Working Group on Planetary System

Nomenclature has announced that the satellite S/2005 S 1 (cf. IAUC 8524) has been given the designation (and name) Saturn XXXV (Daphnis).

COMET 177P/BARNARD Comet P/1889 M1 = 2006 M3 (cf. IAUC 8726) has been given the

permanent number 177P (cf. MPC 57086).

THE EDGAR WILSON AWARD 2006 The Smithsonian Astrophysical Observatory announces that the 2006

Edgar Wilson Award for the discovery of comets (cf. IAUC 6936, 8554) is being divided among the following three individuals: Charles Wilson Juels (Fountain Hills, AZ, U.S.A.) and Paulo Renato Centeno Holvorcem (Campinas, Brazil) for C/2005 N1; and John Broughton (Reedy Creek, Qld., Australia) for P/2005 T5.

V2576 OPHIUCHI D. K. Lynch, R. J. Rudy, C. C. Venturini, and S. Mazuk, Aerospace

Corporation; R. C. Puetter, University of California at San Diego; and R. B. Perry, Langley Research Center, NASA, report 0.47- to 2.5-μm spectroscopy of V2576 Oph (cf. IAUC 8700, 8710) on June 14.4 UT using the VNIRIS spectrograph on the Lick 3-m telescope. This Fe II nova has broad (2700 km/s), flat-topped emission lines. There is very little reddening [E(B − V) = 0.25] and no evidence of dust in the nova ejecta. The object displays weak coronal lines from [S VIII] and [Si VI], and the unidentified features at 1.11, 1.19, 1.55, and 2.10 μm are present. Magnitudes for V2576 Oph: V = 13.9, J = 11.2, H = 11.7, K = 11.0.

2006 July 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green